# Identification of Dynamic Nonlinear Panel Models

### v2

@@ Abstract
This paper studies identification for a wide range of nonlinear panel data models, including binary choice, ordered repsonse, and other types of limited dependent variable models. Our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of (potentially contemporary) endogeneity. Our identification strategy relies on a partial stationarity condition, which not only allows for an unknown distribution of errors but also for temporal dependencies in errors. We derive partial identification results under flexible model specifications and provide additional support conditions for point identification. We demonstrate the robust finite-sample performance of our approach using Monte Carlo simulations, and apply the approach to analyze the empirical application of income categories using various ordered choice models.
Keywords: Panel Discrete Choice Models; Stationarity; Dynamic Models; Partial Identification; Endogeneity
@@ 1 Introduction
This paper studies semiparametric partial identification of a wide range of panel data models with limited dependent variables, including various discrete choice models, censored dependent variable models, and nonseparable models. In particular, our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of endogenous covariates.
To fix ideas, we start with the following dynamic binary choice model, which is on its own of considerable theoretical and applied interest. Sections 3-6 generalize the approach to other limited dependent variable models. Specifically, consider
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (1)
where $Y_{it}\in\{0,1\}$ denotes a binary outcome variable for individual $i=1,2,...$ and time $t=1,...,T$, while $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$. The objective is to identify the parameter $\theta_{0}:=(\beta_{0}^{\prime},\gamma_{0}^{\prime})^{\prime}$ using a panel of observed variables $(Z_{it},X_{it},Y_{it})_{it}$. We focus on short panels, where the number of time periods $T\geq 2$ is fixed and finite.
@@
$\epsilon_{is}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\ \stackrel{{\scriptstyle d}}{{\sim}}\ \epsilon_{it}\mid Z_{is},Z_{it},X_{is},X_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (2)
@@
This condition is also referred to as “group stationarity” or “group homogeneity” and has also been exploited in studies such as Chernozhukov et al. (2013), Shi, Shum, and Song (2018) and Pakes and Porter (2022). Condition (2) does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows dependence between the fixed effect $\alpha_{i}$ and the covariates $(Z_{it},X_{it})$. However, condition (2) does impose substantial restriction on the dependence between $(Z_{it},X_{it})$ and the time-varying error term $\epsilon_{it}$: it effectively requires that all covariates in $(Z_{it},X_{it})$ are exogeneous with respect to the time varying error $\epsilon_{it}$.222 For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=X_{it}^{\prime}\eta$, then the conditional distributions of $\epsilon_{it}$ and $\epsilon_{is}$ cannot be the same as long as $X_{it}^{\prime}\eta\neq X_{is}^{\prime}\eta$, so condition (2) fails in general.
@@
In many economic applications, certain components of the observable covariates, namely $X_{it}$, may exhibit endogeneity. For example, in a dynamic setting where $X_{it}$ includes the lagged outcome variable $Y_{i,t-1}$, then endogeneity of $Y_{i,t-1}$ with respect to $\epsilon_{i,t-1}$ arises immediately. For another example, if $X_{it}$ includes “price” or other variables that may be endogenously chosen by economic agents, then the exogeneity restriction imposed by condition (2) will again fail to hold.
@@
We propose in this paper a general identification approach for various nonlinear panel models in the presence of endogenous regressors, using the following “partial stationarity” condition, which can be viewed as a weaker verision of the stationarity condition (2) above:
$\epsilon_{is}\mid Z_{is},Z_{it},\alpha_{i}\ \stackrel{{\scriptstyle d}}{{\sim}}\ \epsilon_{it}\mid Z_{is},Z_{it},\alpha_{i}\quad\text{for any}\ s,t\leq T.$ (3)
@@
Our partial stationary condition (3), as its name suggests, only requires that the errors are stationary across time periods conditional on the realizations of a subvector of the covariates (i.e., the exogenous covariates denoted by $Z$) while allowing the remaining covariates (denoted by $X$) to be endogenous in arbitrary manners. In short, condition (3) imposes exogeneity conditions only on exogenous covariates. Alternatively, we can interpret condition (3) as an assumption of the existence of some covariates being exogenous. [^3] Condition (3) also accommodates the standard stationarity assumption conditional on all covariates.
@@
We describe how to exploit the partial stationarity condition (3) to derive the identified set on the model parameters $\theta_{0}$ through conditional moment inequalities, which take the form of lower and upper bounds for the conditional distribution $\epsilon_{it}+\alpha_{i}\mid Z_{is},Z_{it}$, solely as functions of observed variables and the model parameters $\theta_{0}$. We show that these bounds must have nonzero intersections over time under the partial stationarity assumption, thereby forming identifying restrictions for the parameter $\theta_{0}$.
While extensive work exists on nonlinear panel models under stationarity, previous studies typically focus on individual models. The identification strategies are often context specific and the identification results may have various complicated representations, as seen in studies such as Khan, Ponomareva, and Tamer (2023) and Pakes and Porter (2022). Our approach offers the advantage of providing a simple and unified characterization of the identified set for a broad range of static and dynamic panel models, irrespective of the specific types of variables (discrete/continuous outcome and covariates) and forms of endogeneity (lagged/contemporary endogenous regressors). Furthermore, our strategy does not rely on scalar-additive structures that are often imposed in various parametric and semiparametric models, and can be further extended to accommodate nonseparable panel data models.
@@
We demonstrate the sharpness of the identified set for binary choice models when the support of $X_{it}$ is finite. More precisely, we show that, for any $\theta$ that satisfies all the conditional moment inequalities we derived, we can construct an observationally equivalent joint distribution of the observed and unobserved variables in our model. While our main result is about set identification, we also provide sufficient conditions for the point identification of the coefficients $\beta_{0}$ on exogenous covariates (under scale normalization) as well as the signs of the coefficients on endogenous covariates $\gamma_{0}$.
@@
Our identification strategy based on partial stationarity can be applied more broadly beyond the context of dynamic binary choice models. We demonstrate its applicability to other limited dependent variable models, such as ordered response models, multinomial choice models, and censored outcome models. The results of our approach accommodates both static and dynamic settings across all these models. Furthermore, we illustrate the adaptability of our key strategy to nonseparable semiparametric models with endogeneity.
@@
Our paper contributes directly to the line of econometric literature on semiparametric panel discrete choice models. Dating back to Manski (1987), a series of work exploits “full” stationarity conditions for identification, such as Chernozhukov et al. (2013), Khan, Ponomareva, and Tamer (2016), Shi, Shum, and Song (2018), Pakes and Porter (2022), Khan, Ouyang, and Tamer (2021), Khan, Ponomareva, and Tamer (2023), Gao and Li (2020), and Wang (2022). As discussed above, full stationarity conditions given all observable covariates effectively require that all covariates are exogenous with no dynamic effects (i.e., lagged dependent variables). In contrast, we exploit the “partial” stationarity condition, thereby allowing for lagged dependent variables as well as other endogenous covariates.
@@
In the literature on dynamic discrete choice models, our paper is most closely related to Khan, Ponomareva, and Tamer (2023, KPT thereafter), who studies the following dynamic panel binary choice model
$Y_{it}=\mathbbm{1}\{Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (4)
where the one-period lagged dependent variable $Y_{i,t-1}\in\{0,1\}$ serves as the endogenous covariate, and $Z_{it}$ are exogenous covariates. KPT exploits the following assumption:
$\epsilon_{is}\mid Z_{it},Z_{is},\alpha_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid Z_{it},Z_{is},\alpha_{i}\quad\text{for any}\ s,t\leq T.$
@@
which is exactly a “partial stationarity” condition in the specific context of (4), and derives the sharp identified set for $\theta_{0}$ by explicitly enumerating the realizations of the one-period lagged outcome variable $Y_{i,t-1}$ (across two periods $t,s$). In contrast, our model (1), along with the “partial stationarity” condition, is stated with more general specifications of the endogenous covariates $X_{it}$. The covariates $X_{it}$ can include more than one lagged dependent variables (e.g. $Y_{i,t-1},Y_{i,t-2},$) and other endogenous variables (such as “price” if $Y_{it}$ represents the purchase of a particular product), which may be continuously valued. Consequently, our identification strategy is substantially different from that of KPT, and can be applied more broadly to various other dynamic nonlinear panel models. In the specific model (4), we show that the identifying restrictions we derived are equivalent to those derived in KPT and thus both approaches lead to sharp identification. Relatedly, Mbakop (2023) proposes a computational algorithm to derive conditional moment inequalities in a general class of dynamic discrete choice models (potentially with multiple lags). The focus of Mbakop (2023) is on scenarios where the lagged discrete outcome variables are the only endogenous covariates in the model, and the proposed algorithm relies on the discreteness of these variables. Relative to these works, our paper introduces an analytic approach that directly applies to a more general class of dynamic binary choice models, as well as other types of models with continuous limited dependent variables and any number of endogenous covariates, regardless of whether they are discrete or continuous and whether they take the form of lagged outcome variables or not.
@@
Our identification strategy relies on a type of stationarity condition, while alternative approaches utilize other notions of exogeneity. For example, Aristodemou (2021) exploits a type of “full independence” assumption to identify dynamic binary choice models. The “full independence” assumption essentially requires that the time-varying errors from all time periods and the exogenous variables from all time periods are independent (conditional on initial conditions), but does not make intertemporal restrictions on the errors (such as stationarity”. Hence, such ‘full independence” assumption and the partial stationarity assumption in our paper do not nest each other as special cases. Chesher, Rosen, and Zhang (2023) applies the framework of generalized instrumental variables (Chesher and Rosen, 2017) to the context of various dynamic discrete choice models with fixed effects, and utilizes a similar “full independence” assumption (Aristodemou, 2021) for identification. [^4] Our identification strategy shares some conceptual similarity with the idea of generalized instrumental variable (GIV) in Chesher and Rosen (2017), who proposes a general approach for representing the identified set of structural models with endogeneity. Chesher and Rosen (2017), Chesher and Rosen (2020), and Chesher, Rosen, and Zhang (2023) demonstrate how the GIV framework can be applied to various settings, but focus mostly on the use of exclusion restrictions and/or full independence assumptions. In this paper, we neither impose exclusion restrictions nor independence assumptions but instead explore identification under a partial stationarity condition. More differently, some other papers work with sequential exogeneity in various dynamic panel models and provide (non-)identification results under different model restrictions. For example, Shiu and Hu (2013) imposes a high-level invertibility condition along with a restriction that rules out the dependence of covariates on past dependent variables. More recently, Bonhomme, Dano, and Graham (2023) investigates panel binary choice models with a single binary predetermined covariate under sequential exogeneity, whose evolution may depend on the past history of outcome and covariate values. The sequential exogeneity condition considered in these papers and the partial stationarity condition in ours again do not nest each other as special cases: in particular, our current paper accommodates contemporarily endogenous covariates that violate sequential exogeneity. In summary, the key assumptions, identification strategy, and identification results of these studies are substantially different from and thus not directly comparable to those in our current paper.
@@
The rest of the paper is organized as follows. Section 2 studies the sharp identification of binary choice models with endogenous covariates. We provide sufficient conditions to achieve point identification and explore two classic models: static and dynamic models. Sections 3-5 apply the identification approach to ordered response models, multinomial choice models, and models with censored dependent variables. Section 6 broadens the scope to general nonseparable models. Section 7 presents simulation results about the finite-sample performances of our approach and Section 8 explores the empirical application of income categories using various ordered response models. We conclude with Section 9.

### v3

# Identification of Nonlinear Dynamic Panels

@@ Abstract
This paper provides a general identification approach for a wide range of nonlinear panel data models, including binary choice, ordered response, and other types of limited dependent variable models. Our approach accommodates dynamic models with any number of lagged dependent variables as well as other types of (potentially contemporary) endogeneity. Our identification strategy relies on a partial stationarity condition, which not only allows for an unknown distribution of errors but also for temporal dependencies in errors. We derive partial identification results under flexible model specifications and provide additional support conditions for point identification. We demonstrate the robust finite-sample performance of our approach using Monte Carlo simulations, and apply the approach to analyze the empirical application of income categories using various ordered choice models.
Keywords: Dynamic Panel, Discrete Choice, Endogeneity, Stationarity, Partial Identification
@@ 1 Introduction
Panel models have been a crucial tool in economics, as they not only allow for individual heterogeneity but also for investigating dynamic effects over time. However, identifying dynamic panel models can be challenging, especially for models with limited dependent variables. Firstly, we may encounter nonseparable and unknown functional forms, rendering it unable to remove individual fixed effects through simple demeaning. Furthermore, lagged dependent variables in dynamic models introduce additional endogeneity, posing another challenge for identification.
This paper provides a general and unified identification approach for a wide range of panel data models with limited dependent variables, including various discrete (binary, multinomial, and ordered) choice models and censored dependent variable models. In particular, our approach accommodates dynamic models with any number of lagged dependent variables as well as contemporarily endogenous covariates. Moreover, the identification approach does not impose parametric distributions on unobserved heterogeneity, and thus it is robust to distributional specifications.
To fix ideas, we start with the following dynamic binary choice model, which is on its own of considerable theoretical and applied interest. Section 3 generalizes the approach to other limited dependent variable models. Specifically, consider
$Y_{it}=\mathbf{\mathbbm{1}}\{Z_{it}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (1)
where $Y_{it}\in\{0,1\}$ denotes a binary outcome variable for individual $i=1,2,...$ and time $t=1,...,T$, while $Z_{it}\in\mathcal{R}^{d_{z}}$ denotes exogenous covariates, $X_{it}\in\mathcal{R}^{d_{x}}$ denotes potentially endogenous covariates, $\alpha_{i}\in\mathcal{R}$ denotes the unobserved fixed effect for individual $i$, and $\epsilon_{it}$ denotes the unobserved time-varying error term for individual $i$ at time $t$. The objective is to identify the parameter $\theta_{0}:=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$ using a panel of observed variables $\left(Z_{i},X_{i},Y_{i}\right){}_{i=1}^{n}$, where $Z_{i}:=\left(Z_{i1},...,Z_{iT}\right)$, and simiarly $X_{i},Y_{i}$ are vector representations of $X_{it},Y_{it}$. We focus on short panels, where the number of time periods $T\geq 2$ is fixed and finite.
@@
$\epsilon_{is}\,\sim\,\epsilon_{it}\mid\alpha_{i},Z_{i},X_{i}\quad\forall s,t=1,...,T$ (2)
@@
This condition is also referred to as “group stationarity” or “group homogeneity” and has also been exploited in studies such as Chernozhukov, Lee, and Rosen (2013), Shi, Shum, and Song (2018) and Pakes and Porter (2022). [^1] To be precise, condition (2) is often stated in the following weaker “pairwise” version in the literature, $\epsilon_{is}\,\sim\,\epsilon_{it}\mid\alpha_{i},Z_{is},Z_{it},X_{is},X_{it},\ \quad\forall s,t=1,...,T,$ where only covariate realizations from the two periods $\left(s,t\right)$ are conditioned on. However, the difference between condition (2) and the pairwise version above usually only leads to minor adaption of the results in the aforementioned papers (as well as in the current one). See Remark 4 for a follow-up discussion. Condition (2) does not impose parametric restrictions on the distributions of $\epsilon_{it}$ and allows dependence between the fixed effect $\alpha_{i}$ and the covariates $\left(Z_{i},X_{i}\right)$. However, condition (2) does impose substantial restriction on the dependence between $\left(Z_{i},X_{i}\right)$ and the time-varying error term $\epsilon_{it}$: it effectively requires that all covariates in $\left(Z_{i},X_{i}\right)$ are exogeneous with respect to the time varying error $\epsilon_{it}$.222 For instance, suppose $\mathbb{E}[\left.\epsilon_{it}\right|Z_{is},Z_{it},X_{is},X_{it}]=X_{it}^{\prime}\eta$, then the conditional distributions of $\epsilon_{it}$ and $\epsilon_{is}$ cannot be the same as long as $X_{it}^{\prime}\eta\neq X_{is}^{\prime}\eta$, so condition (2) fails in general.
@@
In many economic applications, certain components of the observable covariates, namely $X_{i}$, may exhibit endogeneity. For example, in a dynamic setting where $X_{it}$ includes the lagged outcome variable $Y_{i,t-1}$, then endogeneity of $Y_{i,t-1}$ with respect to $\epsilon_{i,t-1}$ (and all $\epsilon_{i,s}$ with $s\geq t$) arises immediately. For another example, if $X_{it}$ includes “price” or other variables that may be endogenously chosen by economic agents after observing $\epsilon_{it}$, then $X_{it}$ would be correlated with contemporary $\epsilon_{it}$, so the exogeneity restriction imposed by condition (2) will again fail to hold.
@@
In this paper, we instead impose and exploit the following “partial stationarity” condition, which can be viewed as a weaker verision of the condition (2) above:
$\epsilon_{is}\sim\epsilon_{it}\mid\alpha_{i},Z_{i},\quad\forall s,t=1,...,T.$ (3)
@@
Our partial stationary condition (3), as its name suggests, only requires that the errors are stationary conditional on the realizations of a subvector of the covariates (i.e., the exogenous covariates denoted by $Z_{i}$) while allowing the remaining covariates (denoted by $X_{i}$) to be endogenous in arbitrary manners. [^3] Our identification strategy and results can be easily adapted under the alternative “pairwise parital stationarity” condition $\epsilon_{is}\sim\epsilon_{it}\mid\alpha_{i},Z_{is},Z_{it}.$ See Remarks 4 and 5 for follow-up discussions. In short, condition (3) imposes exogeneity conditions only on exogenous covariates. Alternatively, we can interpret condition (3) as an assumption of the existence of some covariates being exogenous. [^4] Condition (3) also accommodates the standard stationarity assumption conditional on all covariates.
@@
We describe how to exploit the partial stationarity condition (3) to derive the identified set on the model parameters $\theta_{0}$ through a class of conditional moment inequalities, which take the form of lower and upper bounds for the conditional distribution $\epsilon_{it}+\alpha_{i}\mid Z_{i}$, solely as functions of observed variables and the model parameters $\theta_{0}$. We show that these bounds must have nonzero intersections over time under the partial stationarity assumption, thereby forming a class of identifying restrictions for the parameter $\theta_{0}$. Conditional on the exogenous covariates $Z_{i}$, our class of inequalities is indexed by a scalar $c\in\mathcal{R}$, which implicitly traces out all possible values that the parametrix index $Z_{it}^{{}^{\prime}}\beta_{0}+X_{it}^{{}^{\prime}}\gamma_{0}$ can take. That said, we show how the effective number of identifying restrictions can be reduced to be finite when $X_{it}$ has finite support, a condition naturally satisfied in the important special case of “$p$-th order autorgressive” dynamic binary choice models, where $X_{it}$ consists of lagged outcome variables $Y_{i,t-1},Y_{i,t-2},...,Y_{i,t-p}$ that are by construction discrete.
@@
We demonstrate the sharpness of the identified set we derived for binary choice models under the discreteness of the endogenous covariates $X_{it}$. More precisely, we show that, for any $\theta$ that satisfies all the conditional moment inequalities we derived, we can construct an observationally equivalent joint distribution of the observed and unobserved variables in our model. Our proof of sharpness consists of two steps: first we demonstrate “per-period” sharpness, and then generalize the result to “all-period” sharpness. A key innovation in our proof technique is using an explicit, simple, and general construction that shows how marginal/aggregate stationarity restrictions and joint choice probability restrictions can be satisfied simultaneously, which might be of independent and wider use. While our main result is about set identification, we also provide sufficient conditions for the point identification of the coefficients on exogenous covariates (under scale normalization) as well as the signs of the coefficients on endogenous covariates.
@@
Our identification strategy based on partial stationarity applies more broadly beyond the context of dynamic binary choice models. In Section 3, we demonstrate its applicability in a general nonseparable semiparametric model under monotonicity, and shows how it can be applied to a range of alternative limited dependent variable models, such as ordered response models, multinomial choice models, and censored outcome models. The results of our approach accommodates both static and dynamic settings across all these models.
While extensive work exists on panel discrete choice or other nonlinear panel data models under stationarity-type conditions, previous studies typically focus on individual models. The identification strategies are often context specific and the identification results may have various complicated representations, as seen in studies such as Khan, Ponomareva, and Tamer (2023) and Pakes and Porter (2022). Our approach offers the advantage of providing a simple and unified characterization of the identified set for a broad range of static and dynamic panel models, irrespective of the specific types of variables (discrete/continuous outcome and covariates) and the specific forms of endogeneity (lagged/contemporary endogenous regressors).
@@
Our paper contributes directly to the line of econometric literature on semiparametric panel discrete choice models. Dating back to Manski (1987), a series of work exploits “full” stationarity conditions for identification, such as Chernozhukov, Lee, and Rosen (2013), Khan, Ponomareva, and Tamer (2016), Shi, Shum, and Song (2018), Pakes and Porter (2022), Khan, Ouyang, and Tamer (2021), Khan, Ponomareva, and Tamer (2023), Gao and Li (2020), and Wang (2022). As discussed above, full stationarity conditions given all observable covariates effectively require that all covariates are exogenous with no dynamic effects (i.e., lagged dependent variables). In contrast, we exploit the “partial” stationarity condition, thereby allowing for lagged dependent variables as well as other endogenous covariates.
@@
In the literature on dynamic discrete choice models, our paper is most closely related to Khan, Ponomareva, and Tamer (2023, KPT thereafter), who studies the following dynamic panel binary choice model:
$Y_{it}=\mathbf{\mathbbm{1}}\{Z_{it}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}+\alpha_{i}+\epsilon_{it}\geq 0\},$ (4)
@@
where the one-period lagged dependent variable $Y_{i,t-1}\in\{0,1\}$ serves as the endogenous covariate, and $Z_{it}$ are exogenous covariates. KPT exactly imposes the “partial stationarity” condition (3) in the specific context of (4), and derives the sharp identified set for $\theta_{0}$ by explicitly enumerating the realizations of the one-period lagged outcome variable $Y_{i,t-1}$ (across two periods $t,s$). In contrast, our model (1), along with the “partial stationarity” condition, is stated with more general specifications of the endogenous covariates $X_{it}$. The covariates $X_{it}$ can include more than one lagged dependent variables (e.g. $Y_{i,t-1},Y_{i,t-2},$) and other endogenous variables (such as “price” if $Y_{it}$ represents the purchase of a particular product), which may be continuously valued. Consequently, our identification strategy is substantially different from that of KPT, and can be applied more broadly to various other dynamic nonlinear panel models. In the specific model (4), we show that the identifying restrictions we derived are equivalent to those derived in KPT and thus both approaches lead to sharp identification. Relatedly, Mbakop (2023) proposes a computational algorithm to derive conditional moment inequalities in a general class of dynamic discrete choice models (potentially with multiple lags). The focus of Mbakop (2023) is on scenarios where the lagged discrete outcome variables are the only endogenous covariates in the model, and the proposed algorithm relies on the discreteness of these variables. Relative to these works, our paper introduces an analytic approach that directly applies to a more general class of dynamic binary choice models, as well as other types of models with continuous limited dependent variables and any number of endogenous covariates, regardless of whether they are discrete or continuous and whether they take the form of lagged outcome variables or not.
@@
Our identification strategy relies on a type of stationarity condition, while alternative approaches utilize other notions of exogeneity. For example, Aristodemou (2021) exploits a type of “full independence” assumption to identify dynamic binary choice models. The “full independence” assumption essentially requires that the time-varying errors from all time periods and the exogenous variables from all time periods are independent (conditional on initial conditions), but does not make intertemporal restrictions on the errors (such as stationarity). Hence, such ‘full independence” assumption and the partial stationarity assumption in our paper do not nest each other as special cases. Chesher, Rosen, and Zhang (2023) applies the framework of generalized instrumental variables (Chesher and Rosen, 2017) to the context of various dynamic discrete choice models with fixed effects, and utilizes a similar “full independence” assumption (Aristodemou, 2021) for identification. [^5] Our identification strategy shares some conceptual similarity with the idea of generalized instrumental variable (GIV) in Chesher and Rosen (2017), who proposes a general approach for representing the identified set of structural models with endogeneity. Chesher and Rosen (2017), Chesher and Rosen (2020), and Chesher, Rosen, and Zhang (2023) demonstrate how the GIV framework can be applied to various settings, but focus mostly on the use of exclusion restrictions and/or full independence assumptions. In this paper, we neither impose exclusion restrictions nor independence assumptions but instead explore identification under a partial stationarity condition. More differently, some other papers work with sequential exogeneity in various dynamic panel models and provide (non-)identification results under different model restrictions. For example, Shiu and Hu (2013) imposes a high-level invertibility condition along with a restriction that rules out the dependence of covariates on past dependent variables. More recently, Bonhomme, Dano, and Graham (2023) investigates panel binary choice models with a single binary predetermined covariate under sequential exogeneity, whose evolution may depend on the past history of outcome and covariate values. The sequential exogeneity condition considered in these papers and the partial stationarity condition in ours again do not nest each other as special cases: in particular, our current paper accommodates contemporarily endogenous covariates that violate sequential exogeneity. In summary, the key assumptions, identification strategy, and identification results of these studies are substantially different from and thus not directly comparable to those in our current paper.
@@
The rest of the paper is organized as follows. Section 2 studies the sharp identification of panel binary choice models with endogenous covariates. Sections 3 demonstrates how our key identification strategy generalizes to a wide range of dynamic nonlinear panel data models, such as ordered response models, multinomial choice models, and censored outcome models. Section 4 presents simulation results about the finite-sample performances of our approach and Section 5 explores the empirical application of income categories using various ordered response models. We conclude with Section 6.